OpenStudy (anonymous):

Which expression is true? :A. 50 kilometers/hour > 139 centimeters/second B. 50 kilometers/hour < 139 centimeters/second C. 50 kilometers/hour = 139 centimeters/second D. cannot be determined

OpenStudy (mathstudent55):

You need to convert km/h to cm/s

OpenStudy (anonymous):

how do i do that?

OpenStudy (mathstudent55):

You need to start with basic unit conversions. Do you know how many cm are in a km? If not, then do you know how many m are in a km, and how many cm are in a m? Also, how many seconds are in an hour? If you don't know that, then how many seconds are in a minute, and how many minutes are in an hour?

OpenStudy (anonymous):

Kilometer = 1000 meters 1 meter = 100 centimeters 1 hour = 60 minutes 1 minute = 60 seconds ..... is it A

OpenStudy (mathstudent55):

Ok, your conversions are correct. Now we convert: \(50 \dfrac{km}{h} \times \dfrac{1000 ~m}{1 ~km} \times \dfrac{100~cm}{1~m} \times \dfrac{1~h}{60~mn} \times \dfrac{1~mn}{60~s} \) Notice that all units cancel out except cm/s. Just multiply out the numbers.

OpenStudy (anonymous):

i think i did it right so it should be A right?

OpenStudy (mathstudent55):

Yes. I got 50 km/h = 1388.888... cm/s, so 50 km/h > 139 cm/s.